OpenStudy (anonymous):

-6a + 3 = -3(2a - 1) 1. a = 5 2. a = 10 3. no solution 4. identity

pooja195 (pooja195):

−6a+3=−3(2a−1) −6a+3=(−3)(2a)+(−3)(−1)(Distribute) −6a+3=−6a+3 Add 6a to both sides. −6a+3+6a=−6a+3+6a 3=3 Subtract 3 from both sides. 3−3=3−3 0=0 So what would the answer be?

OpenStudy (anonymous):

No solution